OpenStudy (anonymous):

Write the rule for the following arithmetic sequence: 1, -6, -13, -20,…

OpenStudy (danjs):

arithmetic has a constant change between each term, here it is -7...

OpenStudy (danjs):

linear

OpenStudy (danjs):

you notice for any term, the next term is always 7 less than the previous

OpenStudy (danjs):

then the first term is 1, so you want to start at that value...

OpenStudy (danjs):

a1, a2, a3, a4, …, an ; where a2 = a1 + d, a3 = a2 + d, and so on. The general form will be an = a1 + (n-1)*d

OpenStudy (danjs):

For any term n, the value is that, the first value + (n-1) times the constant value between terms, here d=-7

OpenStudy (danjs):

a1=1

OpenStudy (danjs):

you get all that or?

OpenStudy (anonymous):

yes i did thank you.